OpenStudy (anonymous):

help how do you solve y=x=5 y=3x

RhondaSommer (rhondasommer):

can help

OpenStudy (mathmale):

The first equation is not correct. Fix it, please.

OpenStudy (anonymous):

sry y=x+5

RhondaSommer (rhondasommer):

okay so subtract 5 from both sides

RhondaSommer (rhondasommer):

of that equation

OpenStudy (anonymous):

ok

RhondaSommer (rhondasommer):

what did you get?

OpenStudy (anonymous):

-5+3x=x

RhondaSommer (rhondasommer):

it shoule be y-5=x

RhondaSommer (rhondasommer):

should*

RhondaSommer (rhondasommer):